OpenStudy (anonymous):

integral [e^(1/x^7)]/x^8 dx any hints for me please?

zepdrix (zepdrix):

\[\Large \int\limits e^{\color{orangered}{1/x^7}}\left(\frac{1}{x^8}dx\right)\] U substitution, let the orange part be your u. Your du should look very similar to the stuff in the brackets. It might take a little bit of wiggling around to make them match though.

OpenStudy (anonymous):

ah thank you, i got it now ^^